# Cohort-Based Active Modality Acquisition

### v2

@@ 5 Evaluation
Figure 2: End-to-end architectures to determine the scores for different AFs in our proposed CAMA setting. (A) Vanilla late fusion (LF) architecture of a model $f$ that can handle missing data modalities by masking. If all modalities are available, the model creates scores $s_{i}^{\text{full}}$, otherwise scores $s_{i}^{\text{avail}}$ are created. (B) Possible architecture for training (left) and inference (right) with a LF model $f$ and a generative model $f_{\text{imp}}$ to create scores $s_{i}^{\text{imp}}$ for the imputation-based AFs.
To evaluate these strategies in practice, we require architectures that produce the necessary scores. The oracle, upper-bound, baseline, and random AFs can be evaluated using a vanilla discriminative late fusion (LF) model (Figure˜2 A), as they operate on true labels ${\textnormal{y}}_{i}$ and true scores ${\textnormal{s}}_{i}^{\textnormal{full}}$ and ${\textnormal{s}}_{i}^{\textnormal{avail}}$ (Section˜3). In contrast, our proposed imputation-based AFs are grounded in counterfactual reasoning: They require the model to predict how its output would change if a missing modality were present. This necessitates a more sophisticated architecture that combines the discriminative classifier with a generative component capable of imputing the missing modality (Figure˜2 B).
@@
We implement a multimodal, discriminative, end-to-end LF architecture $f$ combined with generative models $f_{\text{imp}}$ for imputing the missing embeddings (Figure˜2 B). Here, $f$ is composed of modality-specific encoders, , e.g., Vision Transformers (ViTs) (Dosovitskiy et al., 2021) for images and BERT (Devlin et al., 2019) for text, and a classifier for fusing the modality embeddings, i.e., a Transformer encoder (Vaswani et al., 2017). For the generative component $f_{\text{imp}}$, we primarily employ DDPMs that model conditional mappings between latent spaces, i.e., $p({\bm{z}}_{\text{missing}}\,|\,{\bm{z}}_{\text{available}})$, to sample embeddings for the absent modality (Wang et al., 2023) (Figure˜4 in Appendix˜D). Other models are also suitable, e.g., we implement Beta-Conditional-VAEs (BC-VAEs) (Higgins et al., 2017) (Appendix˜F) instead of DDPMs to illustrate architectural trade-offs. During training, the imputed modality embeddings (latents) produced by $f_{\text{imp}}$ are not passed to the classification head, even when modalities are missing (Figure˜2 B, left). Instead, the discriminative components (encoders and head) are trained only on actually available modalities; imputed latents are masked out via the attention mask in the fusion module. The generative and discriminative parts are trained jointly, but the generative outputs do not directly influence the classifier during training beyond the shared encoders being updated by the classification loss. During inference, the reconstructed predictions from $f_{\text{imp}}$ are also sent to the classifier (Figure˜2 B, right), because the imputation-based AFs expect counterfactuals to predict the imputed scores $s_{i}^{\text{imp}}$. The final loss function for the architecture is defined as
$\mathcal{L}=\lambda_{1}\cdot\mathcal{L}_{\text{CE}}+\sum_{i=1}^{P}\lambda_{2}\cdot\mathcal{L}_{\text{D}_{i}}$ (6)
@@
with loss weightings $\lambda$, the classification loss $\mathcal{L}_{\text{CE}}$, the (generative) DDPM losses $\mathcal{L}_{\text{D}_{i}}$, and the total number of modalities $P$. We find that setting $\lambda_{1}=1.0$ is important for downstream performance. For model training, we use the ScheduleFree optimizer (Defazio et al., 2024) with hyperparameters determined through sweeps. As a detail, we find three architectural decisions essential: (a) applying Layer Normalization (Ba et al., 2016) at the end of each modality’s encoder to stabilize the DDPMs operating between latent spaces, (b) calibrating the model with Label Smoothing (Szegedy et al., 2016) to produce less overconfident and better-calibrated probability distributions, and (c) implementing the score models of the DDPMs as Diffusion Transformers (DiTs) (Peebles & Xie, 2023) for improved resource efficiency. Regarding missing modalities, we do not pre-train on all available data modalities, in contrast to Wang et al. (2023). We use a predefined, seed-dependent missing-modality mask to control data modality leakage during training unlike batch-dependent masks, which eventually reveal all modalities for every sample across numerous epochs. Further details in Appendix˜D.
@@
We evaluate the setting of CAMA on four real-world multimodal datasets: UKBB (Sudlow et al., 2015), MIMIC Symile (Saporta et al., 2024), MIMIC HAIM (Soenksen et al., 2022a; b), and MOSEI (Zadeh et al., 2018), which cover diverse domains such as healthcare and emotion recognition. For the publicly available datasets, missing modalities are synthetically created, whereas for UKBB they are an inherent characteristic. We design the datasets for binary classification, resulting in ten binary targets for the MIMIC datasets and one binary target for MOSEI and UKBB. While MOSEI is already class-balanced (Zadeh et al., 2018), HAIM and Symile exhibit significant class imbalance (Soenksen et al., 2022a; Zadeh et al., 2018). To address this, we employ random oversampling during training, which we find essential for the effective operation of AFs. Importantly, during testing we retain the original imbalanced distributions, and no class-balancing steps are applied to UKBB. We highlight UKBB as the most challenging dataset to demonstrate that CAMA scales to a broad multimodal range and large-scale cohorts with approximately 100,000 samples and 15 modalities. In this setting, we focus on acquiring the exceptionally costly proteomics data for predicting the onset of systemic lupus erythematosus (SLE), which has been shown to benefit from proteomics combined with other clinical data (Yang et al., 2025). Additional details are provided in Appendix˜E.
Figure 3: (a) AUROC curves for several AFs on the MOSEI dataset (Zadeh et al., 2018) at an acquisition budget of $25\%$ of the dataset size. (b) Acquisition performance of the best-performing AF from (a), visualizing the gain achieved during the progressive acquisition of modalities as the cohort transitions from unimodal scores towards full multimodality. Notably, the oracle AF can exceed the full multimodal cohort’s AUROC at certain fractions of acquired modalities before subsequently declining towards it again.
@@
For datasets with at least three modalities, we apply five-fold cross-validation. Due to initially noisy results for MIMIC HAIM, we increase the number of folds to ten. For each sample in the test set, the initial score ${\textnormal{s}}_{i}^{\textnormal{avail}}$ is established by randomly assigning a subset of available modalities ${\mathcal{P}}_{i}^{\textnormal{avail}}$. This procedure is repeated over several runs for robustness. In each run, every sample is stochastically assigned a new subset ${\mathcal{P}}_{i}^{\textnormal{avail}}$. Performance metrics are averaged across these independent runs to ensure our evaluation is robust to any single random assignment of patient data. Acquisition is simulated by incrementally increasing the budget $\beta$. We focus on tasks where the multimodal model demonstrates a performance improvement over the unimodal baseline. For certain prediction tasks, a simpler unimodal model can outperform a more complex multimodal one, potentially due to the introduction of noisy or conflicting signals. In such cases, the final multimodal performance falls below the unimodal baseline, resulting in a negative normalized area of gain, indicating that acquisition was detrimental. To ensure a meaningful evaluation, we exclude any tasks exhibiting this negative gain from the analysis at the split level. For each budget, the top-ranked samples in $\mathcal{S}$} are considered acquired, and their logits are updated from ${\textnormal{s}}_{i}^{\textnormal{avail}}$ to ${\textnormal{s}}_{i}^{\textnormal{full}}$. Final reported results are aggregated across all cross-validation splits, combinations of missing and available modalities, and random runs to ensure robustness of the evaluation.
@@ 6 Results
Our empirical evaluation confirms the effectiveness of CAMA. We benchmark our imputation-based strategies against oracles, upper-bound heuristics, and baselines across multiple datasets. Full results are aggregated in Tables˜3, 5 and 5 by averaging over permutations of missing input modalities and multiple random instantiations for each missingness configuration. As expected, oracle strategies serve as an upper bound and consistently achieve the highest performance. Surprisingly, oracle gains can exceed the value of one, as a strategic mix of unimodal and multimodal samples can outperform a purely multimodal cohort. To benchmark the acquisition logic itself, we use label-agnostic upper-bound heuristics that access full scores ${\textnormal{s}}_{i}^{\textnormal{full}}$. Among these, strategies based on KL-Divergence and rank change perform well, indicating that prioritizing large predictive shifts or cohort reordering is an effective heuristic in this setting. Our main approach for handling the CAMA setting comprises imputation-based strategies that leverage a generative model $f_{\text{imp}}$ to predict counterfactual outcomes. The imputation-based KL-Divergence strategy consistently and significantly outperforms all other non-oracle methods. This AF effectively identifies samples predicted to have the largest shift in their class probability distribution (Figure˜3). In contrast, imputation-based strategies relying on rank change, final uncertainty, or final probability are considerably weaker, suggesting that quantifying the change in prediction is more effective than estimating the final state. While our primary results with respect to imputation-based AFs use DDPMs, a BC-VAE variant offers significantly faster inference for a minor trade-off in performance (Tables˜6 and F). The relative performance ranking of these strategies is largely consistent across all datasets, including the large-scale UKBB cohort with approximately 100,000 samples and 15 modalities. This confirms the robustness and scalability of our framework in a challenging setting. In summary, our results affirm the superiority of the imputation-based KL-Divergence strategy, which achieved substantial and reliable gains over all baselines and heuristic methods. Additional results in Appendices˜G, H and I.
7 Discussion
@@
We introduce CAMA to address the challenge of strategic data acquisition under budget constraints. Our experiments consistently demonstrate that imputation-based AFs provide a robust and effective solution. In the following, we discuss the key implications. The ability of oracles to yield gains exceeding that of a model using complete multimodal data for all samples (Figure˜3 (b)), suggests that an underlying predictive model can achieve better global performance with a strategic curation of samples, rather than applying all modalities across the cohort. This likely occurs because additional modalities may introduce variance, redundancy, or conflicting information that imperfect models cannot optimally reconcile. The oracles circumvent this by selecting only additional modalities beneficial to the global metric. To our surprise, the imputation-based KL-Divergence AF can slightly outperform the corresponding upper-bound heuristic (Table˜5). Conversely, the substantial performance gap between the rank-change heuristic and its imputation-based counterpart suggests that global, rank-based metrics may be particularly vulnerable to imputation noise. While the KL-Divergence AF demonstrated strong performance, not all imputation-based AFs consistently outperformed simpler strategies across all datasets or endpoints (Appendices˜H and I). This indicates that optimal CAMA AFs can be context-dependent and that effectiveness hinges on how imputations are leveraged rather than on imputation quality alone. We show CAMA s robustness to imputation errors since $f_{\text{imp}}$ models a distribution of plausible outcomes rather than aiming for a single reconstruction. By averaging the expected impact across this distribution, the acquisition decision becomes less sensitive to uncertainty. Additionally, the primary KL-Divergence AF is resilient to noise, as it prioritizes samples expected to cause a large predictive shift, effectively ignoring minor imputation errors. Taken together, CAMA is not only practical for constrained settings, but also reveals insights into multimodal behavior. The successful KL-Divergence strategy and the surprising oracle performance underscore that the value of an additional modality is not absolute but highly contextual. The most effective AFs are not those that simply predict an outcome, but estimate the magnitude of the predictive shift.
@@ 8 Conclusion and Future Work
We introduce CAMA, a novel setting addressing the real-world challenge of optimizing global discriminative performance through strategic test-time acquisition of additional modalities under resource constraints. Our evaluation across multiple multimodal datasets shows that imputation-based AFs can effectively guide resource allocation under cohort-level constraints. The generally consistent relative ordering of AFs across diverse datasets and the low variance in overall results lend confidence to the robustness of our core findings. In settings such as healthcare, strategic allocation of costly or invasive diagnostic procedures is essential, and our approach offers a promising direction for these applications. Future work includes extending CAMA to multi-class problems or regression tasks, exploring additional imputation techniques, directly optimizing cohort-level metrics, and dynamically selecting which modality to acquire.

### v3

@@ 5 Evaluation
Figure 2: End-to-end architectures to determine the scores for different AFs in our proposed CAMA setting. (A) Vanilla late fusion (LF) architecture of a model $f$ that can handle missing data modalities by masking. The model creates scores $s_{i}^{\text{avail}}$ given the available modalities. (B) Architecture for training (left) and inference (right) with a LF model $f$ and a generative model $f_{\text{imp}}$ to create scores $s_{i}^{\text{imp}}$ for the imputation-based AFs.
To evaluate these strategies in practice, we require architectures that produce the necessary scores. The oracle, upper-bound, baseline, and random AFs can be evaluated using a vanilla discriminative late fusion (LF) model (Figure˜2 A), as they operate on true labels ${\textnormal{y}}_{i}$ and true scores ${\textnormal{s}}_{i}^{\textnormal{acquired}}$ and ${\textnormal{s}}_{i}^{\textnormal{avail}}$ (Section˜3). In contrast, our proposed imputation-based AFs are grounded in counterfactual reasoning: They require the model to predict how its output would change if a missing modality were present. This necessitates a more sophisticated architecture that combines the discriminative classifier with a generative component capable of imputing the missing modality (Figure˜2 B).
@@
First, we implement a multimodal architecture consisting of modality-specific encoders $f_{E}^{(m)}:\mathcal{X}^{(m)}\to\mathbb{R}^{d}$ and a fusion classifier $f_{C}$ (Figure˜2 A). The encoders map raw inputs for a sample $i$ and modality $m$ to latent embeddings $z_{i}^{(m)}=f_{E}^{(m)}(x_{i}^{(m)})$, e.g., with a Vision Transformer (ViT) (dosovitskiy_image_2021) for images and BERT (devlin_bert_2019) for text. The discriminative classifier $f_{C}$ aggregates these embeddings to produce logits $s_{i}^{\text{avail}}$ depending on the availability of the raw inputs, i.e., with a Transformer encoder (vaswani_attention_2017). Second, for our generative AFs, we incorporate generative modules additionally to the discriminative late fusion (Figure˜2 B) (wang_incomplete_2023). The generative modules $f_{\text{imp}}$ are parameterized as Diffusion Transformers (DiTs) (peebles_scalable_2023) or Beta-Conditional-VAEs (BC-VAEs) (higgins_beta-vae_2017) trading off performance vs. efficiency. The generative modules $f_{\text{imp}}$ are trained to approximate the conditional distribution $p_{\theta}(z^{(k)}|\{z^{(m)}\}_{m\in\mathcal{P}^{\text{avail}}_{i}})$ for each target modality $k$. The generative loss for sample $i$ is:
$\mathcal{L}_{G_{i}}=\sum_{k\in\mathcal{P}^{\text{avail}}_{i}}\mathcal{L}_{\text{gen}}\!\left(z_{i}^{(k)};\,\{z_{i}^{(m)}\}_{m\in\mathcal{P}^{\text{avail}}_{i}}\right)$ (6)
where $\mathcal{L}_{\text{gen}}$ is a variational bound on the negative conditional log-likelihood: for DDPMs, this corresponds to the denoising objective (ho_denoising_2020); for BC-VAEs, this is the negative conditional evidence lower bound (ELBO) (higgins_beta-vae_2017). For the discriminative task, the loss for sample $i$ is defined as binary cross entropy loss with the label $y$ and the predicted probability $p$:
$\mathcal{L}_{\text{CE}_{i}}=-\left[y_{i}\log(p_{i})+(1-y_{i})\log(1-p_{i})\right].$ (7)
The final loss function for the whole architecture is defined as the combination of both loss terms:
$\mathcal{L}=\lambda_{1}\mathcal{L}_{\text{CE}}+\lambda_{2}\mathcal{L}_{\text{G}}$ (8)
Figure 3: (a) AUROC curves for several AFs on the MOSEI dataset (zadeh_multimodal_2018) at an acquisition budget of $25\%$ of the dataset size. (b) Acquisition performance of the best-performing AF from (a), visualizing the gain achieved during the progressive acquisition of modalities as the cohort transitions from pre-acquisition scores towards post-acquisition. Notably, the oracle AF can exceed the post-acquisition cohort’s AUROC at certain fractions of acquired modalities before subsequently declining towards it again.
@@
with loss weightings $\lambda$ for which we find $\lambda_{1}=\lambda_{2}=1$ is important for downstream performance. During inference, the classifier $f_{C}$ also uses samples of $p(z^{(k)}|\{z^{(m)}\}_{m\in\mathcal{P}^{\text{avail}}_{i}})$ for missing modalities to create the scores $s_{i}^{\text{imp}}$ needed for our generative AFs (Figure˜2 B, right). During training, samples from $p(z^{(k)}|\{z^{(m)}\}_{m\in\mathcal{P}^{\text{avail}}_{i}})$ are not passed to $f_{C}$, even when modalities are missing (Figure˜2 B, left). Instead, the discriminative components ($f_{E}^{(m)}$ and $f_{C}$) are trained only on available modalities via attention masks. This means that the generative and discriminative parts are trained jointly, but the generative outputs do not directly influence the classifier during training beyond the shared encoders being updated by the classification loss. For model training, we use the ScheduleFree optimizer (defazio_road_2024) with hyperparameters determined through sweeps. We find the following architectural decisions essential, which are ablated in Table˜6: (a) applying Layer Normalization (ba_layer_2016) at the end of each modality’s encoder to stabilize the DDPMs operating between latent spaces, (b) calibrating the model with label smoothing (szegedy_rethinking_2016) to produce less overconfident and better-calibrated probability distributions, (c) decoupling the generative modules from the classifier during training and (d) class balancing the training dataset as detailed in the next paragraph. Regarding missing modalities, we do not pre-train on all available data modalities, in contrast to wang_incomplete_2023. We use a predefined, seed-dependent missing-modality mask to control data modality leakage during training unlike batch-dependent masks, which eventually reveal all modalities for every sample across numerous epochs. Further details in Appendix˜D.
@@
We evaluate the setting of CAMA on four real-world multimodal datasets: UKBB (sudlow_uk_2015), MIMIC Symile (saporta_contrasting_2024), MIMIC HAIM (soenksen_integrated_2022; soenksen_code_2022), and MOSEI (zadeh_multimodal_2018), which cover diverse domains such as healthcare and emotion recognition. For the publicly available datasets, missing modalities are synthetically created, whereas for UKBB they are an inherent characteristic. We design the datasets for binary classification, resulting in ten binary targets for the MIMIC datasets and one binary target for MOSEI and UKBB. While MOSEI is already class-balanced (zadeh_multimodal_2018), HAIM and Symile exhibit significant class imbalance (soenksen_integrated_2022; zadeh_multimodal_2018). To address this, we employ random oversampling during training, which we find essential for the effective operation of AFs (Table˜6). Importantly, during testing we retain the original imbalanced distributions, and no class-balancing steps are applied to UKBB. We highlight UKBB as the most challenging dataset to demonstrate that CAMA scales to a broad multimodal range and large-scale cohorts with approximately 100,000 samples and 15 modalities. In this setting, we focus on acquiring the exceptionally costly proteomics data for predicting the onset of systemic lupus erythematosus (SLE), which has been shown to benefit from proteomics combined with other clinical data (yang_plasma_2025). Details in Appendix˜E.
@@
For datasets with at least three modalities, we apply five-fold cross-validation. Due to initially noisy results for MIMIC HAIM, we increase the number of folds to ten. For each sample in the test set, the initial score ${\textnormal{s}}_{i}^{\textnormal{avail}}$ is established by randomly assigning a subset of available modalities ${\mathcal{P}}_{i}^{\textnormal{avail}}$. This procedure is repeated over several runs for robustness. In each run, every sample is stochastically assigned a new subset ${\mathcal{P}}_{i}^{\textnormal{avail}}$. Performance metrics are averaged across these independent runs to ensure our evaluation is robust to any single random assignment of patient data. Acquisition is simulated by incrementally increasing the budget $\beta$. We focus on tasks where the post-acquisition model demonstrates a performance improvement over the pre-acquisition baseline. For certain prediction tasks, a simpler pre-acquisition model can outperform a more complex post-acquisition one, potentially due to the introduction of noisy or conflicting signals. In such cases, the final post-acquisition performance falls below the pre-acquisition baseline, resulting in a negative normalized area of gain, indicating that acquisition was detrimental. To ensure a meaningful evaluation, we exclude any tasks exhibiting this negative gain from the analysis at the split level. For each budget, the top-ranked samples in $\mathcal{S}$} are considered acquired, and their logits are updated from ${\textnormal{s}}_{i}^{\textnormal{avail}}$ to ${\textnormal{s}}_{i}^{\textnormal{acquired}}$. Final reported results are aggregated across all cross-validation splits, combinations of missing and available modalities, and random runs to ensure robustness.
@@ 6 Results
Our empirical evaluation confirms the effectiveness of CAMA. We benchmark our imputation-based strategies against oracles, upper-bound heuristics, and baselines across multiple datasets. Full results are aggregated in Tables˜3, 5 and 5 by averaging over permutations of missing input modalities
Table 6: Cross-validated ablation of the proposed model adjustments on the Symile dataset, exemplary for the mean across all endpoints with the expected KL-Divergence and acquisitions by AUROC.
Ablation $\text{G}_{\text{full}}$ $\uparrow$
KL-Divergence (w.r.t. Table˜3) 0.833
w/o Layer Norm 0.772
w/o label smoothing 0.746
w/o decoupled data flow 0.599
w/o balanced train set 0.568
and multiple random instantiations for each missingness configuration. As expected, oracle strategies serve as an upper bound and consistently achieve the highest performance.
@@
Surprisingly, oracle gains can exceed the value of one, as a strategic mix of pre-acquisition and post-acquisition samples can outperform a purely post-acquisition cohort. To benchmark the acquisition logic itself, we use label-agnostic upper-bound heuristics that access acquired scores ${\textnormal{s}}_{i}^{\textnormal{acquired}}$. Among these, strategies based on KL-Divergence and rank change perform well, indicating that prioritizing large predictive shifts or cohort reordering is an effective heuristic in this setting. Our main approach for handling the CAMA setting comprises imputation-based strategies that leverage a generative model $f_{\text{imp}}$ to predict counterfactual outcomes. The imputation-based KL-Divergence strategy consistently and significantly outperforms all other non-oracle methods. This AF effectively identifies samples predicted to have the largest shift in their class probability distribution (Figure˜3). In contrast, imputation-based strategies relying on rank change, final uncertainty, or final probability are considerably weaker, suggesting that quantifying the change in prediction is more effective than estimating the final state. While our primary results with respect to imputation-based AFs use DDPMs, a BC-VAE variant offers significantly faster inference for a minor trade-off in performance (Tables˜7 and F). The relative performance ranking of these strategies is largely consistent across all datasets, including the large-scale UKBB cohort with approximately 100,000 samples and 15 modalities. This confirms the robustness and scalability of our framework in a challenging setting. In summary, our results affirm the superiority of the imputation-based KL-Divergence strategy, which achieved substantial and reliable gains over all baselines and heuristic methods. Additional results in Appendices˜G, H and I.
7 Discussion
@@
We introduce CAMA to address the challenge of strategic data acquisition under budget constraints. Our experiments consistently demonstrate that imputation-based AFs provide a robust and effective solution. In the following, we discuss the key implications. The ability of oracles to yield gains exceeding that of a model using post-acquisition data for all samples (Figure˜3 (b)), suggests that an underlying predictive model can achieve better global performance with a strategic curation of samples, rather than applying all modalities across the cohort. This likely occurs because additional modalities may introduce variance, redundancy, or conflicting information that imperfect models cannot optimally reconcile. The oracles circumvent this by selecting only additional modalities beneficial to the global metric. To our surprise, the imputation-based KL-Divergence AF can slightly outperform the corresponding upper-bound heuristic (Table˜5). Conversely, the substantial performance gap between the rank-change heuristic and its imputation-based counterpart suggests that global, rank-based metrics may be particularly vulnerable to imputation noise. While the KL-Divergence AF demonstrated strong performance, not all imputation-based AFs consistently outperformed simpler strategies across all datasets or endpoints (Appendices˜H and I). This indicates that optimal CAMA AFs can be context-dependent and that effectiveness hinges on how imputations are leveraged rather than on imputation quality alone. Regarding the impact of imputation quality, it is important to note that we impute latent embeddings optimized for the discriminative task rather than raw data. Consequently, standard generative metrics (like Fréchet inception distance (FID)) are not applicable for comparing imputation quality across different generative models since every generative model influences the encoders latent spaces indirectly. While we observe that utilizing stronger generative models, e.g., DDPMs, results in higher acquisition performance compared to weaker models, e.g., VAEs, our findings indicate that the generative imputation quality is not the only factor. As detailed in Table˜6, the coherence of the overall architecture design, i.e., ensuring the classifier is robust to the distribution of imputed latents, is equally critical for effective acquisition. We show CAMA s robustness to imputation errors since $f_{\text{imp}}$ models a distribution of plausible outcomes rather than aiming for a single reconstruction. By averaging the expected impact across this distribution, the acquisition decision becomes less sensitive to uncertainty. Additionally, the primary KL-Divergence AF is resilient to noise, as it prioritizes samples expected to cause a large predictive shift, effectively ignoring minor imputation errors. Taken together, CAMA is not only practical for constrained settings, but also reveals insights into post-acquisition behavior. The successful KL-Divergence strategy and the surprising oracle performance underscore that the value of an additional modality is not absolute but highly contextual. The most effective AFs are not those that simply predict an outcome, but estimate the magnitude of the predictive shift.
@@ 8 Conclusion and Future Work
We introduce CAMA, a novel setting addressing the real-world challenge of optimizing global discriminative performance through strategic test-time acquisition of an additional modality under resource constraints. Our evaluation across multiple multimodal datasets shows that imputation-based AFs can effectively guide resource allocation under cohort-level constraints. The generally consistent relative ordering of AFs across diverse datasets and the low variance in overall results lend confidence to the robustness of our core findings. In settings such as healthcare, strategic allocation of costly or invasive diagnostic procedures is essential, and our approach offers a promising direction for these applications. Future work includes extending CAMA to multi-class problems or regression tasks, exploring additional imputation techniques, directly optimizing cohort-level metrics, and dynamically selecting which modality to acquire instead of pre-selecting one.